# Monotonic Path-Specific Effects: Application to Estimating Educational Returns

### v1

@@ 3 Semiparametric, EIF-Based Estimation
Let $O=(X,A,\overline{Z}_{K},\overline{M}_{K},Y)$ denote the observed data, and $\mathcal{P}$ a nonparametric model over $O$ wherein all laws satisfy the positivity assumption described in Section 2. Before proceeding, I define the following auxiliary functions, as introduced in Section 2: $\psi_{km_{k}}\triangleq\mathbb{E}[Y(\overline{1}_{k},m_{k})]$ and $\phi_{k}\triangleq\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]$, for all $k\in[K],M_{0}\triangleq A$. Using the identification results given in Section 2, $\psi_{km_{k}}$ can be written in terms of expectations of observed data:
$\psi_{km_{k}}=\mathbb{E}_{X}\mathbb{E}_{Z_{1}|X,1}\ldots\mathbb{E}_{Z_{k}|X,\overline{Z}_{k-1},\overline{1}_{k}}\mathbb{E}[Y|X,\overline{Z}_{k},\overline{1}_{k},m_{k}].$
@@
$\mu_{km_{k}}^{k}\left(X,\bar{Z}_{k}\right)$ $\triangleq\mathbb{E}\left[Y\mid X,\bar{Z}_{k},\overline{1}_{k},m_{k}\right],$
$\mu_{jm_{k}}^{k}\left(X,\bar{Z}_{j}\right)$ $\triangleq\mathbb{E}\left[\mu_{j+1m_{k}}^{k}\left(X,\bar{Z}_{j+1}\right)\mid X,\bar{Z}_{j},\overline{1}_{j+1}\right]\forall j\in[k-1].$
Further, let $\pi_{km_{k}}\big{(}X,\overline{Z}_{k}\big{)}\triangleq\mathrm{Pr}[M_{k}=m_{k}\mid X,\overline{Z}_{k},\overline{1}_{k}]\forall k\in[K]$, and $\pi_{01}\big{(}X\big{)}\triangleq\text{Pr}[A=1|X]$. The efficient influence function (EIF) of $\psi_{km_{k}}$ is closely related to the EIF for the g-formula, and can be written as
$\psi_{km_{k}}(O)=\sum_{j=0}^{k+1}\varphi_{j}(O),$
@@
A similar EIF-based estimator can be used for $\phi_{k}$ to estimate the $\pi_{k}$ terms in Equation 5. This estimator is based on the following nuisance functions for estimation (see Supplementary Material I for details):
@@
Next, following Kennedy (2022, p. 15), let $\mathbb{IF}:\Psi\rightarrow L_{2}(\mathbb{P})$ denote the operator mapping the functionals $\{\Delta_{k},\pi_{k},\eta_{k}\}:\mathcal{P}\rightarrow\mathbb{R},$ $\forall\in[K]$ to their respective influence functions under the nonparametric model $\mathcal{P}$. Because the $(\Delta_{k},\tau_{k}$) components of the decomposition are linear in $\psi_{km_{k}}$, by linearity of the EIF, $(\mathbb{IF}(\Delta_{k}),\mathbb{IF}(\tau_{k})$) can be expressed as linear combinations of $\varphi_{km_{k}}(O)$. In particular, $\mathbb{IF}(\tau_{k})=\varphi_{(k+1)1}(O)-\varphi_{k,0}(O)$ and $\mathbb{IF}(\Delta_{k})=\varphi_{(k+1)0}(O)-\varphi_{k,0}(O)$. The EIFs of $\eta_{k}$ and $\theta_{k}$, $\forall k\in[K]$ under $\mathcal{P}$ are derived as in Proposition 1:
Proposition 1.
@@
for $k\in\{1,\dots K\}$, with $\theta_{0}=\Delta_{0}$, and where $\mathbb{RIF}(\phi)=\mathbb{IF}(\phi)+\phi,$ denotes the recentered EIF of a parameter (about the truth). Their corresponding EIF-based estimators are (see Supplementary Material I for derivations):
@@
where $\widehat{\mathbb{RIF}}(\phi)=\widehat{\mathbb{IF}}(\phi)+\phi,$ and $\widehat{\mathbb{IF}}(\phi)$ denotes the influence function of a parameter evaluated at estimates of its component nuisance functions (see Supplementary Material I for derivations).
When machine learning estimators are used to compute the nuisance functions, in order to ensure the convergence rates outlined in Proposition 2 below, one could assume Donsker-type conditions for the nuisance function estimators, which restricts the set of estimators available to use. Alternatively, to expand the class of estimators that can be used for estimating the nuisance functions, sample-splitting can be used. In particular, Chernozhukov et al. (2017) suggest a “cross-fitting” procedure, which comprises the following steps: (1) Randomly split data into $J$ folds: $\{S_{1},...S_{J}\}$; (2) For each fold $S_{j}$, use the remaining (j1j-1) folds (training sample) to fit a flexible machine-learning model for each of the nuisance functions involved in the estimating equations; (3) For each observation in $j$ (estimation sample), use estimates of the above models to construct a set of estimated RIF functions for $\Delta_{k}\forall k\in\{0,\dots,K-1\}$, and for $(\pi_{k},\tau_{k},\eta_{k},\theta_{k})\forall k\in[K]$; (4) Compute an estimate of the decomposition components by averaging the estimated RIF functions across all subsamples $S_{1}$ through $S_{J}$. When all nuisance functions are estimated via data-adaptive methods and cross-fitting, the semiparametric efficiency of $\theta_{k}^{\mathrm{rEIF}}$ is given in the following proposition:
Proposition 2.
Semiparametric efficiency. Under Assumption 5, and under suitable regularity conditions (e.g. Chernozhukov et al., 2018), then $\hat{\theta}_{k}^{\text{eif}}$ is semiparametric efficient if $\sum_{j=k}^{k+1}\bigg{[}\sum_{l=0}^{j}R_{n}\big{(}\hat{\pi}_{l1})R_{n}\big{(}\hat{\mu}_{l0}^{j})\bigg{]}+\sum_{j=0}^{k-1}\bigg{[}R_{n}\big{(}\hat{\pi}_{j1})R_{n}\big{(}\hat{\mu}_{j0}^{k-1})+R_{n}\big{(}\hat{\pi}_{j1})R_{n}\big{(}\hat{\mu}_{j1}^{k-1})\bigg{]}+\sum_{j=0}^{k}\bigg{[}\sum_{l=0}^{j}R_{n}\big{(}\hat{\pi}_{l1})R_{n}\big{(}\hat{\gamma}_{l}^{j})\bigg{]}=o(n^{-1/2})$, where $R_{n}\big{(}\cdot\big{)}$ denotes a mapping from a nuisance function to its $L_{2}(P)$ convergence rate, and where $\hat{\mu}_{l0}^{K+1}\triangleq\hat{\mu}_{l1}^{K}$.
@@
When data-adaptive methods are used to estimate the nuisance functions, inference on all components can be conducted via the variance of the empirical analog of the EIF, i.e. $\mathbb{P}_{n}[\big{(}\hat{\psi}_{km_{k}}^{\text{EIF}}\big{)}^{2}]/n$. For example, inference on $\tau_{1}$ can be conducted by estimating $\mathbb{P}_{n}[\big{(}\hat{\psi}_{11}^{\text{EIF}}-\hat{\psi}_{10}^{\text{EIF}}\big{)}^{2}]/n$.
@@ 4 Empirical Analysis
I construct four types of variables: educational transitions, adult earnings, a set of confounders for the effect of high school graduation on subsequent transitions and earnings, and a single set of intermediate confounders for the effect of college completion on subsequent transitions and earnings. My educational transition variables contain a binary treatment denoting whether a respondent had graduated high school by age 22, and three binary mediators denoting whether the respondent had attended a 4-year college by age 22, whether the respondent had received a BA degree by age 29, and whether the respondent had enrolled in a graduate level program by age 29, respectively. I assume that all individuals who make a given educational transition have made all previous educational transitions. Thus, by construction, my coding strategy disallows for cases which violate the monotonicity assumption. [^3] Assuming away cases in which an individual makes a particular educational transition without having made all previous transitions serves as a reasonable approximation to reality. Among the set of individuals who have non-missing earnings information in the NLSY97 (i.e., those who comprise my analytic sample), 94% of individuals observed to attend graduate school by age 29 also completed a BA by age 29; 93% of respondents who completed a BA by age 29 had attended a 4-year college by age 22 (6% of those who completed a BA by age 29 first attended a 4-year college between ages 23 and 26 inclusive), and 99% of respondents who attended a 4-year college by age 22 had also completed high school. My outcome of interest is logged average annual earnings at ages 32-36, which I define to be the (logged) average of a respondent’s self-reported wage, salary income, and business income. Earnings are adjusted for inflation to 2023 dollars using the personal consumption expenditures (PCE) index. After dropping respondents with missing earnings information, I accommodate those with zero earnings by adding a small constant of \$1,000\\$1,000 to observed earnings (though in Supplementary Material F, I replicate my main analyses under alternative definitions of earnings).
@@
A large proportion (just under $50\%$) of respondents are missing information on covariates $X$ and $Z$. For my main analyses, I impute missing values on these covariates via multiple imputation to increase efficiency, but in Supplementary Material E, I replicate these analyses restricted to the sample of respondents with complete information. This exercise produces substantively similar results (for covariate means for each of these analytic samples, see Supplementary Material D). After constructing the analytical sample, I apply both the DML estimator described in Section 3 as well as a parametric, regression-with-residuals (RWR) algorithm (described in Supplementary Material A) to implement the proposed decomposition. For the DML approach, I estimate all nuisance functions, using a super learner composed of the Lasso and random forest and, following Chernozhukov et al. (2017), use five-fold cross-fitting. All weights involved in computing the rEIFs are censored at their $1st$ and $99th$ percentiles. Supplementary Material H gives further details about the particular models required given my assumed data generation process.
Figure 2 shows my estimates of the average total effect (ATE) on log earnings and its direct and continuation components under both the DML and RWR procedures. Both procedures return similar estimates, though deviate in the estimated magnitude of MPSE $\theta_{1}$, and DML estimates come expectedly with a significantly greater amount of precision. The first column shows that the estimated ATE of graduating high school on log earnings under DML (RWR) is $0.67$ ($0.63$), which implies an earnings premium of approximately $96\%$. The next two columns indicate that the vast majority ($69\%$ under DML and $75\%$ under RWR) of the ATE operates directly, i.e. net of college attendance, BA completion and graduate school attendance (MPSE $\theta_{0}$, $A\rightarrow Y$). Specifically, high school graduates who do not proceed to college can be expected to earn on average $0.46$ ($0.47)$ log earnings more than high school non-completers under DML (RWR), an earnings premium of $59\%$.
@@
Second, the fine-grained nature of the MPSE decomposition enables us to trace the continuation effects to their constituent components. In particular, while the direct effect of high school completion is comparable to the direct effect of BA graduation on earnings, suggesting an earnings premium of $59\%$ relative to college attendance without completion, the continuation effect via BA completion that it informs (MPSE $\theta_{2}$, $A\rightarrow M_{1}\to M_{2}\rightarrow Y$) only mediates a small amount of the overall ATE because because $\theta_{2}$ is approximately (plus the small value of $\eta_{2}$) equal to $\Delta_{2}$ scaled by the product $\pi_{1}\cdot\pi_{2}=0.24$. In words, despite the relatively large direct effect of BA completion on earnings, given individuals’ low counterfactual probability of BA completion, this transition is not an important mediating pathway of the total effect of high school completion on earnings. The result is that college attendance without completion mediates high school graduation’s earnings effects as much as BA completion, despite the fact that college attendance without completion yields a much smaller earnings return for high school graduates than BA completion without graduate school attendance does for college enrollees.

### v2

@@ 3 Estimation
Let $O=(X,A,\overline{Z}_{K},\overline{M}_{K},Y)$ denote the observed data, and $\mathcal{P}$ a nonparametric model over $O$ wherein all laws satisfy the positivity assumption described in Section 2. Before proceeding, I define the following auxiliary functions, as introduced in Section 2: $\psi_{km_{k}}\triangleq\mathbb{E}[Y(\overline{1}_{k},m_{k})]$ and $\phi_{k}\triangleq\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]$, for all $k\in[K]$. Using the identification results given in Section 2.4, $\psi_{km_{k}}$ can be written in terms of expectations of observed data:
$\psi_{km_{k}}=\mathbb{E}_{X}\mathbb{E}_{Z_{1}\mid X,1}\ldots\mathbb{E}_{Z_{k}\mid X,\overline{Z}_{k-1},1,\overline{1}_{k-1}}\mathbb{E}[Y\mid X,\overline{Z}_{k},1,\overline{1}_{k-1},m_{k}].$ (11)
@@
$\mu_{km_{k}}^{k}\left(X,\bar{Z}_{k}\right)$ $\triangleq\mathbb{E}\left[Y\mid X,\bar{Z}_{k},1,\overline{1}_{k-1},m_{k}\right],$
$\mu_{jm_{k}}^{k}\left(X,\bar{Z}_{j}\right)$ $\triangleq\mathbb{E}\left[\mu_{j+1m_{k}}^{k}\left(X,\bar{Z}_{j+1}\right)\mid X,\bar{Z}_{j},1,\overline{1}_{j}\right]\forall j\in[k-1].$
This recursive definition of $\mu_{jm_{k}}^{k}\left(X,\bar{Z}_{k}\right)$ is a compact way of expressing the nested expectations in Equation 11. For example, in the case of $k=2$, the recursion becomes
$\mu^{2}_{2m_{2}}(X,Z_{1},Z_{2})=\mathbb{E}[Y\mid X,Z_{1},Z_{2},1,\overline{1}_{1},m_{2}],$
$\mu^{2}_{1m_{2}}(X,Z_{1})=\mathbb{E}_{Z_{2}\mid X,Z_{1},1,\overline{1}_{1}}\!\left[\mu^{2}_{2m_{2}}(X,Z_{1},Z_{2})\right],$
$\mu^{2}_{0m_{2}}(X)=\mathbb{E}_{Z_{1}\mid X,1}\!\left[\mu^{2}_{1m_{2}}(X,Z_{1})\right].$
Thus the counterfactual mean is
$\psi_{2m_{2}}=\mathbb{E}_{X}\!\left[\mu^{2}_{0m_{2}}(X)\right].$
These expressions make explicit that, at each step, we take the expectation of the previous conditional expectation with respect to the distribution of the next confounder, $Z_{j+1}$ given $(X,\bar{Z}_{j},1,\overline{1}_{j})$.
Next, let $\pi_{km_{k}}\big(X,\overline{Z}_{k}\big)\triangleq\mathrm{Pr}[M_{k}=m_{k}\mid X,\overline{Z}_{k},1,\overline{1}_{k-1}]$ $\forall k\in[K]$, and $\pi_{01}\big(X\big)\triangleq\mathrm{Pr}[A=1\mid X]$. The efficient influence function (EIF) of $\psi_{km_{k}}$ is closely related to the EIF for the g-formula, and can be written as
$\psi_{km_{k}}(O)=\sum_{j=0}^{k+1}\varphi_{j}(O),$ (12)
@@
For a proof, see Rotnitzky et al. (2017). The expression above decomposes the efficient influence function for $\psi_{km_{k}}$ into k+2k\!+\!2 components, each corresponding to one layer of the iterated expectation representation in Equation 11. The first term, $\varphi_{0}(O)$, centers the EIF (around zero) by subtracting the target parameter $\psi_{km_{k}}$ from its plug-in estimate $\mu^{k}_{0m_{k}}(X)$. The next $k$ terms, $\varphi_{1}(O),\ldots,\varphi_{k}(O)$, are sequential bias–correction terms. Each $\mu^{k}_{jm_{k}}(X,\bar{Z}_{j})$ is an estimated conditional expectation that appears in the iterated formula for $\psi_{km_{k}}$, and estimation error in these nuisance regressions would normally introduce bias. Each $\varphi_{j}(O)$ therefore takes the form of a weighted residual that subtracts off the discrepancy between two successive levels of the recursion, $\mu^{k}_{jm_{k}}(X,\bar{Z}_{j})-\mu^{k}_{(j-1)m_{k}}(X,\bar{Z}_{j-1})$, and weights it by the inverse probability of the relevant portion of the treatment-mediator history. The last term $\varphi_{k+1}(O)$ plays the same bias-correction role for the outcome regression. It incorporates the residual $Y-\mu^{k}_{km_{k}}(X,\bar{Z}_{k})$ and weights it by the inverse probability of the full treatment–mediator sequence required to identify $\psi_{km_{k}}$. Together, these k+1k\!+\!1 bias–correction terms ensure that the influence function is orthogonal to first–order errors in all nuisance functions, allowing the resulting DML estimator to achieve semiparametric efficiency under standard regularity conditions.
Since the expression above gives the efficient influence function for $\psi_{km_{k}}$ under the nonparametric model $\mathcal{P}$, its variance determines the lowest achievable asymptotic variance for any regular, asymptotically linear estimator of $\psi_{km_{k}}$. Consequently, the semiparametric efficiency bound for any asymptotically linear estimator of $\psi_{km_{k}}$ is $\mathbb{E}\big[(\varphi_{km_{k}}(O))^{2}\big]$.
@@
A similar EIF-based estimator can be used for $\phi_{k}$ to estimate the $\pi_{k}$ terms in Equation 6. This estimator is based on the following nuisance functions for estimation (see Supplementary Material J for details):
@@
Next, following Kennedy (2022, p. 15), let $\mathbb{IF}:\Psi\rightarrow L_{2}(\mathbb{P})$ denote the operator mapping the functionals $\{\Delta_{k},\pi_{k},\eta_{k}\}:\mathcal{P}\rightarrow\mathbb{R},$ $\forall\in[K]$ to their respective influence functions under the nonparametric model $\mathcal{P}$. Because the $(\Delta_{k},\tau_{k}$) components of the decomposition are linear in $\psi_{km_{k}}$, by linearity of the EIF, $(\mathbb{IF}(\Delta_{k}),\mathbb{IF}(\tau_{k})$) can be expressed as linear combinations of $\varphi_{km_{k}}(O)$. In particular, $\mathbb{IF}(\tau_{k})=\varphi_{(k+1)1}(O)-\varphi_{k,0}(O)$ and $\mathbb{IF}(\Delta_{k})=\varphi_{(k+1)0}(O)-\varphi_{k,0}(O)$. The EIFs of $\eta_{k}$ and $\theta_{k}$, $\forall k\in[K]$ under $\mathcal{P}$ are derived as in Theorem 3.1:
Theorem 3.1.
@@
for $k\in\{1,\dots K\}$, with $\theta_{0}=\Delta_{0}$, and where $\mathbb{RIF}(\phi)=\mathbb{IF}(\phi)+\phi,$ denotes the recentered EIF of a parameter (about the truth). Their corresponding EIF-based estimators are (see Supplementary Material J for derivations):
@@
where $\widehat{\mathbb{RIF}}(\phi)=\widehat{\mathbb{IF}}(\phi)+\phi,$ and $\widehat{\mathbb{IF}}(\phi)$ denotes the influence function of a parameter evaluated at estimates of its component nuisance functions (see Supplementary Material J for derivations).
When machine learning estimators are used to compute the nuisance functions, in order to ensure the convergence rates outlined in Theorem 3.2 below, one could assume Donsker-type conditions for the nuisance function estimators, which restricts the set of estimators available to use. Alternatively, to expand the class of estimators that can be used for estimating the nuisance functions, sample-splitting can be used. In particular, Chernozhukov et al. (2017) suggest a “cross-fitting” procedure, which comprises the following steps: (1) Randomly split data into $J$ folds: $\{S_{1},...S_{J}\}$; (2) For each fold $S_{j}$, use the remaining (j1j-1) folds (training sample) to fit a flexible machine-learning model for each of the nuisance functions involved in the estimating equations; (3) For each observation in $j$ (estimation sample), use estimates of the above models to construct a set of estimated RIF functions for $\Delta_{k}\forall k\in\{0,\dots,K-1\}$, and for $(\pi_{k},\tau_{k},\eta_{k},\theta_{k})\forall k\in[K]$; (4) Compute an estimate of the decomposition components by averaging the estimated RIF functions across all subsamples $S_{1}$ through $S_{J}$. When all nuisance functions are estimated via data-adaptive methods and cross-fitting, the semiparametric efficiency of $\theta_{k}^{\mathrm{rEIF}}$ is given in the following Theorem:
Theorem 3.2 (Semiparametric efficiency).
Under Assumption 5, and under suitable regularity conditions (e.g. Chernozhukov et al., 2018), then $\hat{\theta}_{k}^{\text{eif}}$ is semiparametric efficient if $\sum_{j=k}^{k+1}\bigg[\sum_{l=0}^{j}R_{n}\big(\hat{\pi}_{l1})R_{n}\big(\hat{\mu}_{l0}^{j})\bigg]+\sum_{j=0}^{k-1}\bigg[R_{n}\big(\hat{\pi}_{j1})R_{n}\big(\hat{\mu}_{j0}^{k-1})+R_{n}\big(\hat{\pi}_{j1})R_{n}\big(\hat{\mu}_{j1}^{k-1})\bigg]+\sum_{j=0}^{k}\bigg[\sum_{l=0}^{j}R_{n}\big(\hat{\pi}_{l1})R_{n}\big(\hat{\gamma}_{l}^{j})\bigg]=o(n^{-1/2})$, where $R_{n}\big(\cdot\big)$ denotes a mapping from a nuisance function to its $L_{2}(P)$ convergence rate, and where $\hat{\mu}_{l0}^{K+1}\triangleq\hat{\mu}_{l1}^{K}$.
@@
Under the DML estimation procedure, inference for all components of the MPSE decomposition is conducted using the efficient influence functions (EIFs) of the target parameters. When nuisance functions are estimated using cross-fitting and data-adaptive methods, the resulting DML estimators based on the EIFs derived above are asymptotically linear and converge at a $\sqrt{n}$-rate. In particular, under standard regularity conditions (Chernozhukov et al., 2017; Kennedy, 2022), each estimator admits the expansion
$\sqrt{n}(\hat{\theta}-\theta)=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})\big[\phi(O)\big]+o_{p}(1),$
where $\phi(O)$ denotes the corresponding EIF. As a result, $\hat{\theta}$ is asymptotically normal with variance $\mathbb{E}[\phi(O)^{2}]$, which is consistently estimated by the empirical variance of the estimated EIF. Wald-type confidence intervals reported in the main text are constructed using this plug-in variance estimator. For example, inference on $\tau_{1}$ can be conducted by estimating $\mathbb{P}_{n}[\big(\hat{\psi}_{11}^{\text{EIF}}-\hat{\psi}_{10}^{\text{EIF}}\big)^{2}]/n$.
@@ 4 Empirical Analysis
I construct four types of variables: educational transitions, adult earnings, a set of confounders for the effect of high school graduation on subsequent transitions and earnings, and a single set of intermediate confounders for the effect of college completion on subsequent transitions and earnings. My educational transition variables contain a binary treatment denoting whether a respondent had graduated high school by age 22, and three binary mediators denoting whether the respondent had attended a 4-year college by age 22, whether the respondent had received a BA degree by age 29, and whether the respondent had enrolled in a graduate level program by age 29, respectively. I assume that all individuals who make a given educational transition have made all previous educational transitions. Thus, by construction, my coding strategy disallows for cases which violate the monotonicity assumption. [^5] Assuming away cases in which an individual makes a particular educational transition without having made all previous transitions serves as a reasonable approximation to reality. Among the set of individuals who have non-missing earnings information in the NLSY97 (i.e., those who comprise my analytic sample), 94% of individuals observed to attend graduate school by age 29 also completed a BA by age 29; 93% of respondents who completed a BA by age 29 had attended a 4-year college by age 22 (6% of those who completed a BA by age 29 first attended a 4-year college between ages 23 and 26 inclusive), and 99% of respondents who attended a 4-year college by age 22 had also completed high school.
My outcome of interest is logged average annual earnings at ages 32–36. For each respondent and each survey year in this age range, I construct total annual labor-market income by summing wage, salary, and business income, and then compute the (logged) average of these annual totals across this age range. This multi-year average yields a more stable measure of early-adult earnings that smooths over short-term income fluctuations.  [^6] Age 36 is the latest age at which earnings are consistently observed in the NLSY97. Because earnings gains associated with graduate education may materialize later in the life course, I also assess robustness to using a later earnings window. In Supplementary Material G, I re-estimate the full MPSE decomposition using the NLSY79 cohort and measure logged earnings over ages 35–44. The qualitative patterns are highly similar, and the contribution of the graduate-education pathway remains modest, for reasons discussed further in the supplement. Earnings are adjusted for inflation to 2023 dollars using the personal consumption expenditures (PCE) index. After dropping respondents with missing earnings information, I accommodate those with zero earnings by adding a small constant of $\mathdollar 1,000$ to observed earnings (though in Supplementary Material F, I replicate my main analyses under alternative definitions of earnings).
@@
A large proportion (just under $50\%$) of respondents are missing information on covariates $X$ and $Z$. For my main analyses, I impute missing values on these covariates via multiple imputation to increase efficiency, but in Supplementary Material E, I replicate these analyses restricted to the sample of respondents with complete information. This exercise produces substantively similar results (for covariate means for each of these analytic samples, see Supplementary Material D). After constructing the analytical sample, I apply both the DML estimator described in Section 3 as well as a parametric, regression-with-residuals (RWR) algorithm (described in Supplementary Material A) to implement the proposed decomposition. For the DML approach, I estimate all nuisance functions, using a super learner composed of the Lasso and random forest and, following Chernozhukov et al. (2017), use five-fold cross-fitting. All weights involved in computing the rEIFs are censored at their $1st$ and $99th$ percentiles. Supplementary Material I gives further details about the particular models required given my assumed data generation process.
Figure 3 shows my estimates of the average total effect (ATE) on log earnings and its direct and continuation components under both the DML and RWR procedures. Standard errors for RWR estimates are obtained via the non-parametric bootstrap, while standard errors for DML estimates are obtained via the variance of the estimated EIF for each MPSE. Both procedures return similar estimates, though deviate in the estimated magnitude of MPSE $\theta_{1}$, and DML estimates come expectedly with a significantly greater amount of precision. The first column shows that the estimated ATE of graduating high school on log earnings under DML (RWR) is $0.67$ ($0.63$), which implies an earnings premium of approximately $96\%$. The next two columns indicate that the vast majority ($69\%$ under DML and $75\%$ under RWR) of the ATE operates directly, i.e. net of college attendance, BA completion and graduate school attendance (MPSE $\theta_{0}$, $A\rightarrow Y$). Specifically, high school graduates who do not proceed to college can be expected to earn on average $0.46$ ($0.47)$ log earnings more than high school non-completers under DML (RWR), an earnings premium of $59\%$.
@@
Second, the fine-grained nature of the MPSE decomposition enables us to trace the continuation effects to their constituent components. In particular, while the direct effect of high school completion is comparable to the direct effect of BA graduation on earnings, suggesting an earnings premium of $59\%$ relative to college attendance without completion, the continuation effect via BA completion that it informs (MPSE $\theta_{2}$, $A\rightarrow M_{1}\to M_{2}\rightarrow Y$) only mediates a small amount of the overall ATE because $\theta_{2}$ is approximately (plus the small value of $\eta_{2}$) equal to $\Delta_{2}$ scaled by the product $\pi_{1}\cdot\pi_{2}=0.24$. In words, despite the relatively large direct effect of BA completion on earnings, given individuals’ low counterfactual probability of BA completion, this transition is not an important mediating pathway of the total effect of high school completion on earnings. The result is that college attendance without completion mediates high school graduation’s earnings effects as much as BA completion, despite the fact that college attendance without completion yields a much smaller earnings return for high school graduates than BA completion without graduate school attendance does for college enrollees.
@@
5 Conclusion
Several directions for future research follow naturally from the proposed framework. First, in this paper I have considered a decomposition of the average treatment effect for the case of binary monotonic mediators, but many educational processes are more finely graded. Extensions to settings with categorical or multivalued transitions—such as different types of postsecondary institutions, fields of study, or graduate degrees—would further broaden the applicability of the framework. Supplementary Material H outlines one such extension, but generalizing the framework to categorical and continuous mediators remains an open area for future work.
Second, although the MPSE decomposition relaxes cross-world assumptions and permits observed intermediate confounding, it still relies on sequential ignorability. In practice, this assumption may be difficult to satisfy fully in observational settings, particularly when selection into successive educational transitions is driven by unobserved traits such as motivation or ability. Developing alternative identification results that, for example, exploit instrumental variables—long used in the education literature to address selection into schooling—would be a particularly promising extension of the MPSE framework.
Finally, while this paper emphasizes educational attainment, the monotonic structure exploited here arises in many other domains characterized by state-dependent transitions, such as family formation, health progression, or criminal justice contact. This characteristic is particularly salient in demographic phenomena. Certain demographic events are rigid in their monotonicity by definition. For example, researchers may be interested in discerning the degree to which positive effects of marriage on outcomes such as earnings and life satisfaction are undermined by the negative effects of divorce and separation (and, in turn, their mitigation via re-marriage) (Kenney, 2004; Sweeney and Phillips, 2004) – monotonic transitions. Similarly, the effect of parenthood on earnings can be seen as operating directly, through the effect of having a first child net of subsequent children, as well as operating indirectly through the effects of having multiple children. A similar perspective may be taken in a criminal justice context: the total effect of early-stage police contact (such as being searched for contraband) on educational and socio-psychological outcomes can be decomposed into path-specific effects via subsequent arrest and incarceration (Weaver and Lerman, 2010; Kirk and Sampson, 2013; Sugie and Turney, 2017). Applying the MPSE framework to these contexts, and comparing the resulting decompositions across domains, may yield new insights into how life transitions shape later outcomes through both direct and sequential mechanisms.
6 Tables and figures